OpenStudy (anonymous):

Please solve using Cramer's rule Give one value of t for which the system of equations x +3y=ty 3x+y =tx has infinitely many solutions. Please provide me with a worked example ...kind regards. I posted this problem a few days ago and got very helpful solutions using subtraction etc., Someone advised using cramers rule but i got stuck because the RHS of the equation is in the form tx and ty rather than numbers. Please help.

OpenStudy (turingtest):

do you know how to find the determinant of a matrix?

OpenStudy (anonymous):

in order to get infinite many solution you should get determinant =0

OpenStudy (anonymous):

choose such a value of t for which the system determinants =0

OpenStudy (anonymous):

simplifying we can write as x+3y-ty=0 3x-tx+y=0

OpenStudy (anonymous):

@unseenoceans you getting the steps?

OpenStudy (anonymous):

ok take t common in the above two equation to have something simple x+y(3-t)=0 (3-t)x+y=0

OpenStudy (turingtest):

\[D_y=\left|\begin{matrix}1&0\\3-t&0\end{matrix}\right|=0\]hm... this does seem strange to use cramers rule here

OpenStudy (anonymous):

can you find determinant of above system ?

OpenStudy (turingtest):

...but the above is always zero, so that would suggest that the value of t does not matter now I'm confused :P

OpenStudy (anonymous):

@TuringTest i think you missed t with y term in first equation.

OpenStudy (turingtest):

in cramers rule you sub the RHS though

OpenStudy (anonymous):

Yes , I know how to find determinant (ab-cd)

OpenStudy (anonymous):

\[x+3y=ty, 3x+y=tx\]

OpenStudy (anonymous):

this is from the problem sheet

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 3-t \\ 3-t & 1\end{matrix}\right]\]

OpenStudy (anonymous):

find the determinant of above ..let me know .you will get Quadratic equation set the equation equal to zero find the two roots. can you do this?

OpenStudy (anonymous):

let me attempt

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is that 8-6t+t^2

OpenStudy (anonymous):

roots are 2 and 4

OpenStudy (anonymous):

I was advised when I first posted this problem that Cramers rule will be quickest. I am not so sure anymore

OpenStudy (turingtest):

1-(3-t)(3-t)=1-9+6t-t^2=-8+6t-t^2=0 I don't think cramers rule is the fastest, it's easiest by inspection I would say

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

use the above t value and you will have system of infinite solution.

OpenStudy (anonymous):

Apart from the speed of solving the problem, can it be done with cramers rule, Turing?

OpenStudy (anonymous):

you can also use Gauss jordon method to check..reduce the system to echelon form after putting t value and you will get one complete row of zeros

OpenStudy (turingtest):

I am getting confused because sami's way works, but doesn't seem to be using cramers rule completely in the way I am used to

OpenStudy (anonymous):

sami there are 2 t values : (2 and 4) the question asks for one value. Which one is the right one or are they both right?

OpenStudy (turingtest):

it says "give one value" so I guess you could give either

OpenStudy (anonymous):

cramer's rule is good for solving the system that have unique solution .for infinite use Gauss jordon.

OpenStudy (turingtest):

I agree^

OpenStudy (anonymous):

nice tip sami - do i reduce the second row to 0's?

OpenStudy (anonymous):

How do you tell that above has infinite no solutions? DO you determine that from the question or is it because of the RHS : tx, ty?

OpenStudy (anonymous):

the easiest way to check weather the system has infinite solutions is take determinent if its zero you will have infinite solutions lets use t=2 x+3y=2y 3x+y=2x or x+y=0 x+y=0 find determinant

OpenStudy (anonymous):

also look at them these are both the equation of same line. and you may know that there are infinite no of solutions to a line. we get solution when the lines intersects. but (parallel lines intersects at infinity)

OpenStudy (anonymous):

yes, I understand now. Great explanation Sami....much appreciated ...I'll be looking out for you and Turing test next time! Kind regards

OpenStudy (anonymous):

Once again thank you Sami-21 and Turing Test. Goodbye.

OpenStudy (anonymous):

you are Most welcome:) @unseenoceans . i will be glad to help you. that's what i do . i love to help students with their Questions .and make them Understand.